OpenStudy (anonymous):

how do i find the exact value for inverse functions? I do not understand how you know which quadrant it is or anything. Here are some examples: a) Sin^-1 (0) b) Sin^-1 (-1) c) tan^-1 (0) d) sin^-1 (square root 2/2) e) tan^-1 (square root 3) f) cos^-1 (- square root 3/3) either pick one or all of them but step by step would be great thanks!

OpenStudy (tkhunny):

You can think of the other function if you like. It might make more sense: a) Sin^-1 (0) ==> \(\sin(What) = 0\) b) Sin^-1 (-1) ==> \(\sin(What) = -1\) c) tan^-1 (0) ==> \(\tan(What) = 0\) d) sin^-1 (square root 2/2) ==> \(\sin(What) = \sqrt{2}/2\) e) tan^-1 (square root 3) f) ==> \(\tan(What) = \sqrt{3}\) f) cos^-1 (- square root 3/3) ==> \(\cos(What) = 1/\sqrt{3}\) Are you SURE that last one isn't supposed to be TANGENT?

OpenStudy (anonymous):

I'll start..... For sin^-1...your answer MUST be between -pi/2 and pi/2 inclusive. No exceptions. So, in problem (a), sin ^-1 (0) = 0 (which is between -pi/2 and pi/s inclusive). For part (b) sin^-1 (-1) = -pi/2 which is between -pi/2 and pi/2 inclusive. For part (d), sin^-1 (sqrt(2)/2) = pi/4...which is between -pi/2 and pi/2 inclusive. Just to make sure you understand what we are doing, in part (d), when we are finding sin^-1 (sqrt(2)/2)..what we want is to find the angle whose sine has a value of sqrt(2)/2..and that occurs at 45 degrees or pi/4.

OpenStudy (anonymous):

tkhunny yes i am sure the last one is not tangent. I get that Sin(WHAT)=0 but both 0 and 180 deg. have 0 as sin. how do you know which one is the correct answer?

OpenStudy (anonymous):

oh easyaspi thank you! you answered my question!

OpenStudy (tkhunny):

0º and 180º are NOT both solutions. The inverse sine function is defined ONLY on \(-90º \le What \le 90º\). This forces you to discard 180º.